OpenStudy (n0z3kp0n3r):

I've been doing this problem for 2 days and I cant get the answer right. Can someone do it for me. I have like 3 alike this one. I just want to know the steps Lim x->2 ((1/sqx+2)- 1/x)/x-2

satellite73 (satellite73):

\[\frac{1}{\sqrt{x+2}}-\frac{1}{x}\]

satellite73 (satellite73):

like that?

OpenStudy (n0z3kp0n3r):

it's that one

OpenStudy (n0z3kp0n3r):

I know the answer should be 3/16 but I cant get to it

satellite73 (satellite73):

subtract

satellite73 (satellite73):

\[\frac{1}{\sqrt{x+2}}-\frac{1}{x}=\frac{x-\sqrt{x+2}}{x\sqrt{x+2}}\] might be a start

satellite73 (satellite73):

then maybe rationalize the numerator

satellite73 (satellite73):

yeah that will do it

satellite73 (satellite73):

\[\frac{x-\sqrt{x+2}}{x\sqrt{x+2}}\times \frac{x+\sqrt{x+2}}{x+\sqrt{x+2}}\]

satellite73 (satellite73):

leave the denominator in factored form , don't multiply it out then numerator will have a factor of \(x-2\) in it cancel with the \(x-2\) that i have avoided writing

OpenStudy (n0z3kp0n3r):

there is where i get stuck

jimthompson5910 (jim_thompson5910):

@n0z3kp0n3r Using the difference of squares rule, you should get \[\Large \left(x-\sqrt{x+2}\right)\left(x+\sqrt{x+2}\right) = \left(x\right)^2-\left(\sqrt{x+2}\right)^2\] \[\Large \left(x-\sqrt{x+2}\right)\left(x+\sqrt{x+2}\right) = x^2-(x+2)\] \[\Large \left(x-\sqrt{x+2}\right)\left(x+\sqrt{x+2}\right) = x^2-x-2\]

jimthompson5910 (jim_thompson5910):

So \[\Large \frac{x-\sqrt{x+2}}{x\sqrt{x+2}}\] turns into \[\Large \frac{x^2-x-2}{\left(x\sqrt{x+2}\right)\left(x+\sqrt{x+2}\right)}\]

OpenStudy (n0z3kp0n3r):

and I dont know what to do after that

OpenStudy (n0z3kp0n3r):

I still have the (x-2) under that whole equetion

OpenStudy (n0z3kp0n3r):

as a denominator

OpenStudy (mww):

The numerator is actually (x-2)(x+1) so the (x-2) cancels

jimthompson5910 (jim_thompson5910):

so we have this \[\Large \frac{x^2-x-2}{(x-2)\left(x\sqrt{x+2}\right)\left(x+\sqrt{x+2}\right)}\]

OpenStudy (mww):

yes it's easy afterwards when you've cnacelled the (x-2). The final form you can sub in x = 2 directly.

jimthompson5910 (jim_thompson5910):

exactly as @mww states, the numerator x^2-x-2 factors to (x-2)(x+1) and the pesky x-2 term goes away. Once x-2 is gone from the denominator, you can use substitution to evaluate the limit

OpenStudy (n0z3kp0n3r):

I think I got it. Thank you all. <3